Chapter 2, Problem 2/084 Multistep A sprinter practicing for the 200-m dash accelerates uniformly from rest at A and reaches a top speed of 39 km/h at the 68-m mark. He then maintains this speed for the next 66 meters before uniformly slowing to a final speed of 34 km/h at the finish line. Determine the maximum horizontal acceleration which the sprinter experiences during the run. Where does this maximum acceleration value occur? Finish line 39.9 m ▼Part 1 Your answer is partially correct. Try again. Calculate (a) the magnitude of the tangential acceleration near start of the race and (b) the magnitude of tha tangential accelaration near the end of the race. Answers: (a) lal -0.86 (b) la2.39 m/s
Answer #1

tamgemtial acceleration

at = v^2/ (2a) = (39* 5/ 18)^2 / ( 2* 66)

at = 0.8891 m/s^2

========

Comment before rate in case any doubt, will reply for sure... goodluck